# Compatible finite element interpolated neural networks

### v1

@@ 4. Conclusions
To evaluate the performance of the methods, we conducted experiments on both forward and inverse problems. Specifically, in the forward Maxwell problem, we analysed how compatible FEINNs perform across different mesh sizes and trial FE space orders. In order to showcase the applicability of compatible FEINNs to surface PDEs, we also tackled a forward Darcy problem on the unit sphere. In general, we observe that using lower-order test bases allows NNs to beat FEM solutions in terms of $L^{2}$ and $H(\textbf{curl})$ errors for problems with a smooth analytic solution, achieving over two orders of magnitude improvement over FE solutions. In addition, interpolated NNs match the FE solutions in all cases. Preconditioning (i.e., using dual residual norms in the loss function) improves the stability and efficiency of the training process. Our findings also suggest that the strong PDE residual with NNs effectively guides mesh refinement during adaptive training, resulting in NNs that are more accurate than FEM solutions. For inverse problems, we compare the compatible FEINN method with the adjoint NN method for the inverse Maxwell problem with partial or noisy observations. Compatible FEINNs are more accurate than adjoint NNs in most cases, do not require explicit regularisation and are robust against random noise in the observations.

### v2

@@ 3.3.2. Noisy observations
3.3.3. Boundary observations
We conclude the inverse experiments section by attacking a boundary value inverse Maxwell problem. This scenario is particularly challenging as state observations are available only around the boundary, while the state in the interior domain is unknown. To add more complexity, we assume that the unknown coefficient $\kappa$ is discontinuous across the domain $\Omega=[0,1]^{2}$. The analytical solutions are as follows:
$\mathbf{u}(x,y)=\begin{bmatrix}\cos(\pi x)\cos(y)\\ \sin(x)\sin(\pi y)\end{bmatrix},\quad\kappa(x,y)=\begin{cases}1,&\text{ if }y>2x\\ 10,&\text{ if }y\leq 2x\\ \end{cases}.$
The curl of the analytic state and the discontinuous $\kappa$ are depicted in Fig. 14(a) and Fig. 14(d), respectively. A total of $4\times 70$ boundary observations are generated by evaluating the true state along the interior side of the boundary. Their locations are marked as purple dots in Fig. 14(a).
((a)) ×$\boldsymbol{\nabla}\times\mathbf{u}$
((b)) ×$\boldsymbol{\nabla}\times\mathbf{u}^{id}$
((c)) |×$|\boldsymbol{\nabla}\times(\mathbf{u}^{id}-\mathbf{u})|$
((d)) $\kappa$
((e)) $\kappa^{id}$
((f)) $|\kappa^{id}-\kappa|$
Figure 14. Illustration of analytical and NN solutions and the errors for the inverse Maxwell problem with boundary observations. The purple dots in (a) represent the observed data. See caption of Fig. 11 for details on the information being displayed in this figure.
In this experiment, we focus exclusively on the compatible FEINN approach. Since the Dirichlet boundary condition is presented as boundary observations, the interpolation space for $\mathbf{u}_{\mathcal{N}}$ is taken as $U^{1}_{h}$ rather than $U^{1}_{h,0}$ in the previous experiments. To reflect this change in the FE space, the loss (20) is modified as follows:
$\mathscr{L}(\boldsymbol{\kappa}_{\mathcal{N}},\mathbf{u}_{\mathcal{N}})\doteq\left\lVert\mathbf{d}-\mathcal{D}_{h}(\pi_{h}^{1}(\mathbf{u}_{\mathcal{N}}))\right\rVert_{\ell^{2}}+\alpha\left\lVert\tilde{\mathcal{R}}_{h}(\pi^{0}_{h}(\boldsymbol{\kappa}_{\mathcal{N}}),\pi^{1}_{h}(\mathbf{u}_{\mathcal{N}}))\right\rVert_{{V^{1}}^{\prime}},$
where ~$\tilde{\mathcal{R}}(\kappa,\mathbf{u},\mathbf{v})=a(\kappa,\mathbf{u},\mathbf{v})-\ell(\mathbf{v})-\int_{\partial\Omega}(\mathbf{v}\times\mathbf{n})\cdot(\boldsymbol{\nabla}\times\mathbf{u})$ is the modified PDE residual. The boundary integral term in $\tilde{\mathcal{R}}$ arises from integration by parts and the fact that $\mathbf{v}\times\mathbf{n}$ is not necessarily zero on $\partial\Omega$.
The NN structures for $\mathbf{u}_{\mathcal{N}}$ and $\kappa_{\mathcal{N}}$ remain the same as in the noisy observations experiment. The training iterations and penalty coefficients are set to $[150,50,3\times 400]$ and $[0.001,0.003,0.009]$, respectively. The mesh consist of $100\times 100$ squares, and the interpolation space orders are consistent with those in the earlier inverse experiments. We present the trained NN solutions and their errors in the last two columns of Fig. 14. Despite having state observations only around the boundary, the curl of the NN state solution produced by our method are visually indistinguishable from the curl of the true state, as evident from the comparison of Fig. 14(b) and Fig. 14(a). Moreover, the identified NN coefficient as shown in Fig. 14(e) successfully captures the jumps of the true coefficient. The pointwise error of the coefficient, depicted in Fig. 14(f), is small in most regions of the domain, with slightly higher errors near the discontinuity.
Fig. 15 illustrates the training error histories for the compatible FEINN method. All relative errors decrease steadily during training and converge to low levels. Additionally, after a few hundred iterations, $\mathbf{u}_{\mathcal{N}}$ errors consistently fall below those of its interpolation counterpart in both $L^{2}$ and $H(\textbf{curl})$ norms. Overall, the training curves confirm the reliability of our 3-step training strategy and the potential for $\mathbf{u}_{\mathcal{N}}$ to achieve higher accuracy than its FE interpolation when trained with boundary observations.
Figure 15. Relative errors during training by the compatible FEINN method for the inverse Maxwell problem with boundary observations.
Similar to previous experiments, we conduct 100 trials with different NN initial parameters to assess our method’s robustness in the boundary observations case. The resulting box plots for the relative errors are presented in Fig. 16. For the state, the interpolated NNs are highly reliable and accurate, with line-like error boxes positioned at 1% relative errors in both $L^{2}$ and $H(\textbf{curl})$ norms. Although non-interpolated NN solutions result in slightly more scattered boxes, they demonstrate greater accuracy than their their interpolation counterparts. This is reflected in the upper whisker lines for the $L^{2}$ and $H(\textbf{curl})$ relative errors, which are below 0.2% and 0.3%, respectively. Regarding the coefficient, the NNs achieve a 1% error reduction compared to their interpolations, with most $\kappa_{\mathcal{N}}$ errors around 3%. These coefficient error boxes are nearly flat, with an interquartile range of less than 1%. These results highlight our method’s robustness in recovering both the state and discontinuous coefficient of the inverse Maxwell problem using only boundary observations.
Figure 16. Box plots from 100 experiments for the inverse Maxwell problem with boundary observations.
@@ 4. Conclusions
To evaluate the performance of the methods, we conducted experiments on both forward and inverse problems. Specifically, in the forward Maxwell problem, we analysed how compatible FEINNs perform across different mesh sizes and trial FE space orders. To showcase the applicability of compatible FEINNs to surface PDEs, we tackled a forward Darcy problem on the unit sphere. In general, we observe that using lower-order test bases allows NNs to beat FEM solutions in terms of $L^{2}$ and $H(\textbf{curl})$ errors for problems with a smooth analytic solution, achieving over two orders of magnitude improvement over FE solutions. In addition, interpolated NNs match the FE solutions in all cases. Preconditioning of NNs (i.e., using dual residual norms in the loss function) improves the stability and efficiency of the training process. Our findings also suggest that the strong PDE residual with NNs effectively guides mesh refinement during adaptive training, resulting in NNs that are more accurate than or comparable to FEM solutions. For inverse problems, we compare the compatible FEINN method with the adjoint NN method for the inverse Maxwell problem with partial or noisy observations. Compatible FEINNs are more accurate than adjoint NNs in most cases, do not require explicit regularisation and are robust against random noise in the observations. Besides, the FEINN method also achieves high accuracy in both state and coefficient for the inverse problem with boundary observations.
@@ Acknowledgments
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.